OpenStudy (anonymous):

solve the system using the elimination method. 2x+5y=13 6x+2y=-13 below heres my solution

OpenStudy (anonymous):

\[x=-5/2y+13/2, y=4\]

OpenStudy (anonymous):

i dont get why i can't get x..

OpenStudy (anonymous):

\[ \ \ \ 2x+5y=13\\ \ \ \ 6x+2y=-13\\ \text{First, let's multiply the first equation by -3.}\\ \ \ \ -6x-15y=-39\\ \text{Now let's combine the two equations.}\\ \ \ \ -13y=-52\\ \ \ \ y=4\\ \text{Now we can use }y\text{ to find }x.\\ \ \ \ 2x+5y=13\\ \ \ \ 2x=13-5(4)\\ \ \ \ \ \ \ \ =13-20\\\ \ \ \ \ \ \ =-7\\ \ \ \ x=-\frac72 \]

OpenStudy (anonymous):

\[\text{You were on the right track, you just hadn't reduced the expression yet...}\\\text{plug in what you found for }y.\]

OpenStudy (anonymous):

do i plug in the original equation or revised one?

OpenStudy (anonymous):

Any, because all are equations... just pick one which looks like it requires the least amount of work.

OpenStudy (anonymous):

alright. so the solution is (-7/2, 4) ya? :)

OpenStudy (anonymous):

do we have to write it as decimal? (-3.5, 4) ?

OpenStudy (anonymous):

Indeed -- it's the point where the lines intersect. Since the lines all meet at a common point here, you can derive y from x using any of the equations. I don't know about whether you need to write it as a fraction or decimal but they're equal.